# NJUST-KMG at TRAC-2024 Tasks 1 and 2: Offline Harm Potential Identification

### v1

@@ 1. Introduction
The TRAC-2024 Offline Harm Potential Identification task is a critical effort aimed at addressing the pressing issues regarding the impact of online content once taken into a real-world, offline context, broadly the task is to predict whether a specific post is likely to initiate, incite or further exaggerate an offline harm event (viz. riots, mob lynching, murder, rape, etc). With the exponential growth of digital platforms, monitoring the diverse and multilingual content becomes paramount to prevent detrimental consequences in social interactions and individual well-being. This task emphasizes the challenging aspect of understanding nuanced implications embedded within conversations in various Indian languages, highlighting the urgency in developing sophisticated models that can navigate the intricacies of linguistic and cultural nuances. The importance of this task is well-articulated in the task overview paper by Sharma et al. (2023), which delineates the framework and significance of identifying content with potential offline harm effectively.
Our system leverages the synergy of advanced pretrained models Devlin et al. (2018) with the progressive concept of contrastive learning Chen et al. (2020). We harness the rich representations learned by models trained on extensive corpuses and tailor these to our specific context through meticulous fine-tuning. By integrating contrastive learning, we enhance the model’s ability to discern subtleties within the dataset’s multilingual content, crafting a more robust system against the diversity of languages and semantic complexities. The ensemble strategy Dietterich (2000) employed at the testing phase not only solidifies the individual strengths of diverse models but also ensures our system’s resilience and generalization across different data points.

### v2

@@ 1. Introduction
The TRAC-2024 Offline Harm Potential Identification task is a critical effort aimed at addressing the pressing issues Yang et al. (2023) regarding the impact of online content once taken into a real-world, offline context, broadly the task is to predict whether a specific post is likely to initiate, incite or further exaggerate an offline harm event (viz. riots, mob lynching, murder, rape, etc). With the exponential growth of digital platforms, monitoring the diverse and multilingual content becomes paramount to prevent detrimental consequences in social interactions and individual well-being. This task emphasizes the challenging aspect of understanding nuanced implications embedded within conversations in various Indian languages, highlighting the urgency in developing sophisticated models that can navigate the intricacies of linguistic and cultural nuances.
Our system leverages the synergy of advanced pretrained models Devlin et al. (2018) with the progressive concept of contrastive learning Chen et al. (2020), which have extensive applications in various fields such as multi-modal learning Yang et al. (2019a, b), continual learning, semi-supervised learning, etc. We harness the rich representationsYang et al. (2024) learned by models trained on extensive corpuses and tailor these to our specific context through meticulous fine-tuning. By integrating contrastive learning, we enhance the model’s ability to discern subtleties within the dataset’s multilingual content, crafting a more robust system against the diversity of languages and semantic complexities Huo et al. (2018). The ensemble strategy Dietterich (2000) employed at the testing phase not only solidifies the individual strengths of diverse models but also ensures our system’s resilience and generalization across different data points.
@@ 6. Bibliographical References
- Huo et al. (2018) Xuan Huo, Yang Yang, Ming Li, and De-Chuan Zhan. 2018. Learning semantic features for software defect prediction by code comments embedding. In 2018 IEEE international conference on data mining (ICDM), pages 1049–1054. IEEE.
@@
- Yang et al. (2023) Yang Yang, Ran Bao, Weili Guo, De-Chuan Zhan, Yilong Yin, and Jian Yang. 2023. Deep visual-linguistic fusion network considering cross-modal inconsistency for rumor detection. Science China Information Sciences, 66(12):222102.
- Yang et al. (2019a) Yang Yang, Zhao-Yang Fu, De-Chuan Zhan, Zhi-Bin Liu, and Yuan Jiang. 2019a. Semi-supervised multi-modal multi-instance multi-label deep network with optimal transport. IEEE Transactions on Knowledge and Data Engineering, 33(2):696–709.
- Yang et al. (2024) Yang Yang, Jinyi Guo, Guangyu Li, Lanyu Li, Wenjie Li, and Jian Yang. 2024. Alignment efficient image-sentence retrieval considering transferable cross-modal representation learning. Frontiers of Computer Science, 18(1):181335.
@@
- Yang et al. (2019b) Yang Yang, De-Chuan Zhan, Yi-Feng Wu, Zhi-Bin Liu, Hui Xiong, and Yuan Jiang. 2019b. Semi-supervised multi-modal clustering and classification with incomplete modalities. IEEE Transactions on Knowledge and Data Engineering, 33(2):682–695.